You have a  atom (3 protons) which is doubly ionized such that it has just one remaining electron. ()

How much energy, in electron volts, does it take to remove the remaining electron if it starts in the n = 1 orbital?


1
Expert's answer
2021-06-07T09:30:40-0400
"E=E_0\cdot\frac{Z^2}{n^2}=13.6\cdot\frac{3^2}{1^2}=122\text{ eV}."